INFRARED LINE INTENSITIES FOR FORMALDEHYDE FROM SIMULTANEOUS MEASUREMENTS IN THE INFRARED AND FAR INFRARED SPECTRAL RANGES

L. FISSIAUX, Laboratoire Lasers et Spectroscopies, Facultés Universitaires Notre Dame de la Paix, 61 rue de Bruxelles, B-5000 Namur, Belgium; T. FÖLDES, Service de Chimie Quantique et Photophysique, Université Libre de Bruxelles, CP 160/09, 50 avenue F.D. Roosevelt, B-1050 Brussels, Belgium; F. KWABIA TCHANA, Laboratoire Interuniversitaire des Systèmes Atmosphériques, CNRS, Universités de Paris Est Créteil et Paris 7, 61 avenue du Général De Gaulle, F-94010 Créteil cedex, France; L. DAUMONT, Groupe de Spectrométrie Moléculaire et Applications, UMR CNRS 6089, Université de Reims Champagne Ardenne, Campus du Moulin de la Housse, BP 1039, 51067 Reims Cedex 2, France; M. LEPÈRE, Laboratoire Lasers et Spectroscopies, Facultés Universitaires Notre Dame de la Paix, 61 rue de Bruxelles, B-5000 Namur, Belgium; \underline{J. VANDER AUWERA}, Service de Chimie Quantique et Photophysique, Université Libre de Bruxelles, CP 160/09, 50 avenue F.D. Roosevelt, B-1050 Brussels, Belgium.

ranges.^{a}

Formaldehyde (H_{2}CO) is an important intermediate compound in the degradation of the volatile organic compounds (VOCs), including methane, in the terrestrial troposphere. Its observation using optical remote sensing in the infrared range relies on the 3.6 and 5.7 μm absorption bands. Band and individual line intensities have been reported in both With the present work, we aim to also derive infrared line intensities for formaldehyde, however relying on pure rotation line intensities and the known electric dipole moment to determine the particle density. Indeed, because formaldehyde polymerizes or degrades easily, the gas phase may contain polymerization or degradation products. Spectra of H_{2}CO diluted in 10 hPa of N_{2} were therefore simultaneously recorded in the 20 − 60 cm^{−1} and 3.6 μm ranges, respectively using a Bruker IFS125HR Fourier transform spectrometer and a tunable diode laser.

asee A. Perrin, D. Jacquemart, F. Kwabia Tchana, N. Lacome, J. Quant. Spectrosc. Radiat. Transfer 110 (2009) 700-716, and references therein